OpenStudy (anonymous):

64p^2=9 solve the equation

OpenStudy (anonymous):

64p^2-9 =0 (8p+3)(8p-3)=0 p= +-3/8

OpenStudy (anonymous):

Alternatively, take the square root of both sides such that you have \[\pm 8p=3\] Then divide by +/-8 to get the same answer.